# Efficient Multi Subject Visual Reconstruction from fMRI Using Aligned Representations

### v1

@@ 4.2 Results
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
Regular FT 0.112 0.306 75.16% 82.06% 71.99% 72.48% 0.872 0.548
AA 0.115 0.307 76.48% 83.28% 75.37% 73.27% 0.855 0.543
AA + IS 0.113 0.298 77.28% 85.40% 79.63% 75.12% 0.844 0.531
@@
5 Extending to Other Datasets
To demonstrate that our proposed method generalizes across datasets, we conducted experiments on the THINGS dataset. We used a model pre-trained on Subject 1 from the NSD dataset and fine-tuned it on Subject 1 from the THINGS dataset. Since a one-to-one mapping does not exist between the THINGS and NSD datasets, we build an implicit common set by comparing the train images of THINGS with the NSD common set to find a subset of semantically similar image pairs. CLIP is used to identify semantically similar images. In the end,  150 images are chosen as the common set for the THINGS dataset. As shown in Table 4, our AAMax method again outperforms conventional fine-tuning. This result is especially impressive as AAMax was able to semantically align the images even without an explicit alignment. We present additional experiments in Appendix Section B to show that AAMax actually improves alignment in the shared space.
Several factors may explain the observed performance differences between the two datasets. While a plurality of recent fMRI reconstruction reports have focused on analyses of the large and high-quality Natural Scenes Dataset, which measured high-SNR 7 Tesla fMRI data at a  1.8 mm resolution, such data are often out of reach for most neuroscience labs. The modal imaging setup available to many cognitive neuroscience labs is a 3T scanner, such as a Siemens Prisma, which can achieve relatively high-quality images at 2-2.5 mm resolution and reasonable sampling rate (0.5-2 s). The THINGS-fMRI dataset was recently made available, which includes high-quality 3T fMRI data acquired at 2 mm resolution (with multiband acceleration) while participants viewed images drawn from the THINGS object database. While, in principle, this dataset is very similar to the NSD, there are notable differences in SNR (7T vs 3T), stimulus presentation protocol (NSD: 3 s stimulus presentation/1 s ITI; THINGS: 0.5 s stimulus presentation/4 s ITI), stimulus set (NSD: MS-COCO, THINGS: curated object images), and task performed by participants (NSD: recognition memory on each image; THINGS: object or not?). Thus, differences in decoding or reconstruction performance are expected, and not themselves a demonstration that models do not reliably generalize across datasets.
Table 4: Quantitative results on finetuning THINGS-fMRI Subject 1 using NSD Subject 1 as the reference subject. Only the high level pipeline is trained therefore pixelwise metrics are excluded. AAMax outperforms normal finetuning even when an explicit data mapping is not available.
@@ 5 Extending to Other Datasets
THINGS Subj1 Normal FineTune 77.46% 81.07% 71.06% 73.16% 0.911 0.619
THINGS Subj1 AAMax FineTune 78.27% 85.55% 74.10% 81.37% 0.902 0.595
6 Related Work
Recent works have achieved impressive results by mapping fMRI data to latent diffusion model (LDM) spaces  (Takagi & Nishimoto, 2023; Scotti et al., 2023; Lu et al., 2023; Xia et al., 2024), while simultaneously integrating multiple modalities. Despite this progress, these methods have not been thoroughly tested for their generalization performance across a larger population of subjects. In other words, constructing a model for a new subject still requires retraining the model from scratch to ensure optimal performance.
Approaches like Ferrante et al. (2024) have demonstrated cross-subject brain decoding using limited data samples by aligning multiple subjects. Although the authors, like us, leveraged visual stimuli commonly used for multiple subjects, the fidelity of their reconstructed images lags behind current advancements. Scotti et al. (2024), the foundation of our method, produced significant results in image reconstruction under limited data conditions but did not fully exploit the properties of the common representation space. In contrast, our novel approach enables faster convergence and improved performance in a limited data setting by effectively leveraging the structure of the common space.
@@ 7 Conclusions & Discussion
In this work, we present compelling evidence for the existence of a shared visual representation space across subjects and datasets, allowing brain activity from different individuals and experimental settings to be aligned in a common space. Building on this foundation, we introduce Adapter Alignment (AA), a novel and efficient approach for subject-specific alignment in fMRI-based visual reconstruction. Our method, which is more efficient than traditional end-to-end training pipelines, achieves faster convergence and superior performance, particularly in low-data scenarios. Additionally, we propose a greedy image selection algorithm that identifies the most representative subset of images for training new subjects. This approach reduces the amount of data required for fine-tuning by up to 50% without sacrificing performance, addressing the challenge of minimizing data requirements in fMRI-based image reconstruction. Finally, we validate the robustness of our methods by testing them on different data settings.
By leveraging the structure of the common representation space, we demonstrate the feasibility of producing high-quality visual reconstructions with a significantly reduced amount of subject-specific data. Future work could explore the integration of non-linear adapters to enhance concept extraction, potentially improving the performance and flexibility of the model. Another promising direction is to extend our methodology to different pipelines, ensuring its generalizability across various architectures, thereby making the approach architecture-agnostic.
Appendix A Scaling Analysis and Projection to Higher-Dimensional Shared Spaces
We are unable to evaluate adapter alignment on the architectural set up employed by MindEye2 due to the difference in computational resources. While MindEye2 utilizes a 4096-dimensional shared space with BigCLIP and Stable Diffusion XL (SDXL), we operate under compute constraints that limit us to Versatile Diffusion. To address this limitation, we conduct a systematic scaling law analysis to evaluate how shared space dimensionality affects performance and whether adapter alignment (AAMax) continues to be effective at larger scales.
A.1 Empirical Observations
@@
This scaling behavior suggests that our approach remains agnostic to the choice of CLIP space or diffusion model, and the benefits of adapter alignment persist across dimensions.
A.2 Projection to 4096 Dimensions with Stable Diffusion XL
The primary claim we make is that adapter alignment is a training paradigm first and foremost, rather than a specific architectural implementation. Our results demonstrate that adapter alignment is agnostic to both the architectural components (CLIP space, diffusion model) and the scale at which fMRI-to-image reconstruction pipelines are trained. Consequently, it should not be directly compared with works that present large-scale models, which rely on significantly higher compute budgets to achieve their performance benchmarks. Instead, adapter alignment provides a computationally efficient alternative training framework that scales effectively, making fMRI-based reconstruction more accessible even under constrained computational resources.

### v2

@@ 5.4 Generalizing To Different Architectures
MindEye1 Subject 2 Baseline 0.078 0.264 75.33% 84.18% 74.15% 73.66% 0.862 0.524
MindEye1 Subject 2 AAMax 0.103 0.283 81.77% 89.39% 80.31% 78.71% 0.824 0.495
MindEye2 Subject 2 Baseline 0.104 0.354 81.71% 88.06% 78.84% 74.15% 0.845 0.499
MindEye2 Subject 2 AAMax 0.140 0.355 82.22% 89.67% 82.97% 78.79% 0.821 0.483
6 Generalizing to Other Datasets
To demonstrate that our proposed method generalizes across datasets, we conducted experiments on the THINGS dataset. We used a model pre-trained on Subject 1 from the NSD dataset and fine-tuned it on Subject 1 from the THINGS dataset. Since a one-to-one mapping does not exist between the THINGS and NSD datasets, we build an implicit common set by comparing the train images of THINGS with the NSD common set to find a subset of semantically similar image pairs. CLIP is used to identify semantically similar images. In the end,  150 images are chosen as the common set for the THINGS dataset. As shown in Table 4, AAMax again outperforms conventional fine-tuning. This result is especially impressive as AAMax was able to semantically align the images even without an explicit alignment. We present additional experiments in Appendix I to show that AAMax actually improves alignment in the shared space.
@@ 6 Generalizing to Other Datasets
Several factors may explain the observed performance differences between the two datasets. While a plurality of recent fMRI reconstruction reports have focused on analyses of the large and high-quality Natural Scenes Dataset, which measured high-SNR 7 Tesla fMRI data at a 1.8 mm resolution, such data is often out of reach for most neuroscience labs. The modal imaging setup available to many cognitive neuroscience labs is a 3T scanner, such as a Siemens Prisma, which can achieve relatively high-quality images at 2–2.5 mm resolution and reasonable sampling rate (0.5–2 s). The THINGS-fMRI dataset was recently made available, which includes 3T fMRI data acquired at 2 mm resolution (with multiband acceleration) while participants viewed images drawn from the THINGS object database. While, in principle, this dataset is very similar to the NSD, there are notable differences in SNR (7T vs 3T), stimulus presentation protocol (NSD: 3 s stimulus presentation/1 s ITI; THINGS: 0.5 s stimulus presentation/4 s Inter-Trial Interval (ITI)), stimulus set (NSD: MS-COCO, THINGS: curated object images), and task performed by participants (NSD: recognition memory on each image; THINGS: object or not?). Thus, differences in decoding or reconstruction performance are expected, and not themselves a demonstration that models do not reliably generalize across datasets.
Table 4: Quantitative results on fine-tuning THINGS-fMRI Subject 1 using NSD Subject 1 as the reference subject. Only the high level pipeline is trained and therefore, pixelwise metrics are excluded. AAMax outperforms the baseline even when an explicit data mapping is not available.
@@
THINGS Subj1 Baseline Fine-Tune 77.46% 81.07% 71.06% 73.16% 0.911 0.619
THINGS Subj1 AAMax Fine-Tune 78.27% 85.55% 74.10% 81.37% 0.902 0.595
@@ 7 Conclusions & Discussion
In this work, we provide evidence that fMRI signals from different subjects can be structurally aligned within a shared representation space, enabling subject-agnostic modeling of visual brain activity. Building on this, we introduce Adapter Alignment (AA), a novel training paradigm that consistently outperforms traditional end-to-end pipelines in low-data regimes while achieving comparable performance in high-data settings. We further propose a greedy image selection algorithm that identifies representative training images, together reducing the data required for fine-tuning by up to 60% while matching baseline performance, and by 40% while significantly outperforming it. These contributions demonstrate that accurate visual reconstructions can be achieved with substantially less data than previously thought feasible, greatly improving the efficiency of fMRI-based reconstruction.
Our experiments show that AA generalizes across subjects, architectures, and datasets, highlighting its robustness and broad applicability. By leveraging the structure of the shared representation space, we make fMRI-based visual reconstruction more efficient and accessible. Future work could investigate richer adapter designs and extend these alignment strategies to other brain signal modalities, further advancing the goal of generalizable brain-to-image reconstruction.
Appendix Contents
Appendix A Fine-tuning on Different Subjects with MindEye2
The main text presents results on fine-tuning Subject 2 using Subject 1 as reference. Here we present the results for Subject 5 and Subject 7 too with 1 hour of training data. In both cases, AAMax outperforms the baseline.
Table 5: Fine-tuning on Subjects 5 and 7 using MindEye2’s architecture. AAMax outperforms the baseline. The same reference subject checkpoint was used in all the runs and all models were trained with the exact same hyperparameter setup.
Low-Level High-Level
Subject 5 FT 1HR Baseline 0.102 0.358 78.40% 86.61% 79.85% 77.23% 0.839 0.494
Subject 5 FT 1HR AAMax 0.121 0.354 80.68% 88.86% 85.12% 81.89% 0.799 0.469
Subject 7 FT 1HR Baseline 0.103 0.356 78.50% 84.70% 76.81% 74.02% 0.858 0.506
Subject 7 FT 1HR AAMax 0.131 0.390 80.80% 87.86% 80.29% 76.55% 0.836 0.508
Appendix B Generalizing to Different Adapter Architectures
To test the sensitivity of Adapter Alignment to the choice of adapter, we compared several lightweight variants. As shown in Table 6, the choice of adding or removing a non-linearity, or substituting different activation functions (e.g., GELU vs. ReLU), does not significantly affect alignment or reconstruction performance. In practice, simpler architectures tend to perform slightly better, with a single-layer adapter outperforming deeper variants. Our adapter choice in the main experiments was guided by the underlying backbone: we use a nonlinear (linear + GELU) adapter with MindEye1 and retain the linear adapter design of MindEye2 for consistency with the original paper. Overall, these results indicate that the specific adapter design does not affect the validity of Adapter Alignment, and performance is robust across lightweight variants.
Table 6: Ablation on adapter variants for Subject 2 fine-tuned with 1 hour of data under AAMax training. We compare a nonlinear adapter with GELU, a nonlinear adapter with ReLU, and a two-layer linear adapter. Results show that adapter choice has little impact on performance, with all variants yielding similar alignment and reconstruction quality.
Low-Level High-Level
Subject 2 FT 1HR AAMAX linear 0.110 0.277 79.70% 88.61% 80.35% 79.88% 0.815 0.486
Subject 2 FT 1HR AAMAX GELU 0.112 0.259 78.75% 88.04% 78.99% 78.28% 0.824 0.513
Subject 2 FT 1HR AAMAX ReLU 0.101 0.273 80.44% 88.18% 78.72% 78.17% 0.827 0.508
Subject 2 FT 1HR AAMAX 2 linear 0.107 0.268 78.28% 87.12% 78.78% 78.62% 0.827 0.504
Appendix C Generalizing to Different Reference Subjects
We conducted additional experiments (Table 7) to examine how the choice of reference subject influences fine-tuning performance. In these experiments, Subjects 2, 5, and 7 were used as reference subjects, and Subject 1 was fine-tuned with one hour of data. The results show that the reference subject does affect reconstruction quality: Subject 5, which typically achieves the strongest individual reconstructions, yields the best fine-tuning performance, while Subject 7 produces the weakest, consistent with its lower single-subject performance. Nonetheless, Adapter Alignment remains effective in all cases, as AAMax consistently outperforms baseline fine-tuning.
Table 7: Effect of reference subject choice on limited-data fine-tuning. We compare AAMax and normal fine-tuning (160 epochs) when Subjects 2, 5, and 7 are used as reference subjects and Subject 1 is fine-tuned. AAMax consistently outperforms normal fine-tuning regardless of the reference, though overall performance reflects the quality of the chosen reference subject.
Low-Level High-Level
Subject 1 FT From Subject 2 1HR Baseline 0.007 0.268 75.55% 84.00% 75.20% 74.60% 0.862 0.520
Subject 1 FT From Subject 2 1HR AAMAX 0.06 0.277 81.13% 88.99% 79.71% 78.89% 0.824 0.494
Subject 1 FT From Subject 5 1HR Baseline 0.094 0.239 71.54% 80.10% 73.11% 74.02% 0.875 0.537
Subject 1 FT From Subject 5 1HR AAMAX 0.105 0.259 76.77% 86.20% 80.08% 80.14% 0.828 0.496
Subject 1 FT From Subject 7 1HR Baseline 0.010 0.266 74.96% 81.55% 71.71% 73.72% 0.874 0.528
Subject 1 FT From Subject 7 1HR AAMAX 0.012 0.283 79.58% 87.78% 77.48% 77.87% 0.835 0.502
Appendix D Image Selection Algorithm Ablations
D.1 Algorithm Coverage
We next evaluate how well the image selection algorithm covers the common space compared to random sampling. Using the top 20 eigenvectors, we discretize each dimension into bins and, for a subset of $K=250$ images, count the number of empty bins (bins without a selected image). Random selection leaves, on average, around 200 empty bins with a standard deviation of about 10 across 1,000 trials. In contrast, our algorithm consistently reduces this number to 15 empty bins. A permutation test yields a p-value of 0.0002, confirming that the improved coverage is statistically significant. This demonstrates that our algorithm selects images that span the representation space more uniformly along high-variance directions. It is important to note that the number of empty bins in our algorithm remain relatively similar when changing the number of top eigenvectors.
Figure 6: Coverage of the image selection algorithm compared to random sampling. We project embeddings onto the top 20 eigenvectors and count empty bins across 250 selected images. The histogram shows empty-bin counts for 1,000 random subsets, while the dashed line marks our algorithm’s selection. Our method leaves only 15 bins empty, whereas random sampling leaves $\approx 200\pm 10$ on average, indicating significantly more uniform coverage of the representation space ($p=0.0002$).
D.2 Varying the Eigenvector Dimension
We conducted an ablation study on the image selection algorithm to evaluate the effect of varying the number of principal eigenvector dimensions used in the binning process. Specifically, we compared subsets selected using the top 10, 20, 30, and 40 eigenvectors. As shown in Table 8, performance is sensitive to this choice: 20 eigenvectors yield the strongest results, with a secondary improvement at 40, while 10 and 30 dimensions perform relatively worse. Based on these findings, we adopt 20 eigenvectors for all main experiments in the paper.
Table 8: Ablation on the number of eigenvector dimensions used for image selection. We compare random image selection and Adapter Alignment with images selected using our algorithm across different eigenvector settings. Results are reported for Subject 1 as the reference and Subject 2 fine-tuned with 250 images under AAMax training, averaged over 3 runs. The best performance is obtained with 20 eigenvectors, which is used in all main experiments.
Low-Level High-Level
Random Selection 0.112 0.259 78.75% 88.04% 78.99% 78.28% 0.824 0.513
10 EV Dimensions 0.114 0.303 78.99% 86.53% 77.30% 75.23% 0.850 0.541
20 EV Dimensions 0.102 0.267 81.41% 89.55% 80.43% 80.33% 0.823 0.496
30 EV Dimensions 0.112 0.275 79.60% 88.03% 78.58% 76.89% 0.850 0.537
40 EV Dimensions 0.118 0.285 81.78% 89.61% 81.49% 79.55% 0.810 0.491
Appendix E Comparing Different Alignment Training Strategies
To directly compare our approach against related methods, we conducted an experiment fine-tuning Subject 2 with 250 training images under three different training regimes (Table 9). First, we evaluated AA Step 1 only, which corresponds to the setting of Ferrante et al. [Ferrante et al., 2024], where adapters are overfitted to the common space without any subsequent end-to-end training. Second, we implemented a variant inspired by MindBridge [Wang et al., 2024], where the adapter remains unfrozen while the MLP and prior are frozen, and the full end-to-end training is carried out using a combined MSE, CLIP, and prior loss. Finally, we compared these baselines to our full AAMax training regime, which explicitly aligns adapters in the first phase and then fine-tunes the entire pipeline end-to-end.
The results show that AAMax achieves the strongest performance across both low-level and high-level metrics, outperforming both AA Step 1 and the MindBridge-style end-to-end setup. Importantly, this experiment was run under identical architectural settings and without auxiliary enhancements such as additional synthetic data or architectural add-ons. This isolates the effect of the training paradigm itself, demonstrating that when compared purely on training objectives, AAMax provides the most effective strategy for fine-tuning new subjects.
Table 9: Comparison of different training regimes for fine-tuning Subject 2 with 250 training images. We evaluate (1) AAMax (full Adapter Alignment with staged alignment + end-to-end fine-tuning), (2) AA Step 1 only (adapter overfitting without subsequent end-to-end training), and (3) End-to-End with MSE+Prior+CLIP Loss (MLP Frozen) (reset-tuning). AAMax achieves the best performance across both low-level and high-level metrics, demonstrating that explicit staged alignment followed by end-to-end fine-tuning is more effective than either alignment-only or combined end-to-end training when compared under identical architectural conditions.
Low-Level High-Level
AAMAX 0.112 0.259 78.75% 88.04% 78.99% 78.28% 0.824 0.513
AA Step 1 only 0.008 0.240 77.14% 87.04% 77.00% 75.43% 0.857 0.510
End-to-End with MSE+Prior+CLIP Loss (MLP Frozen) 0.007 0.238 76.40% 86.42% 77.72% 75.04% 0.852 0.507
Appendix F Extended Limited Data Experiments
To further validate the efficiency of Adapter Alignment, we repeated our limited-data experiments using Subject 5 as the reference and Subject 1 as the fine-tuning subject. We evaluate fine-tuning with 1, 2, and 4 hours of data and compare baseline end-to-end training to AAMax. As always, AAMax first performs Step 1 adapter alignment by overfitting the adapter, followed by end-to-end fine-tuning. We track performance after 1 epoch of fine-tuning and after 160 epochs of fine-tuning. Even after a single epoch, AAMax consistently outperforms baseline models trained for 160–200 epochs.
We also track performance over fine-tuning epochs for all three data conditions (1, 2, and 4 hours). As shown in Figure 7, Adapter Alignment converges rapidly and maintains a consistent advantage throughout training, while baseline training requires hundreds of epochs to reach comparable accuracy. These results reinforce that AAMax substantially reduces both the data and computational requirements for fine-tuning new subjects.
Table 10: Quantitative results on fine-tuning a new subject with limited data. FT=Fine tuning. The first section shows the result of fine-tuning with 250 (1 hour) images after 1 epoch and after 160 epochs on the baseline and using AAMax. The second section compares performance on 500 (2 hours) images and the third section compares performance on 1000 (4 hours) images. In all cases AAMax, significantly outperforms the baseline. The greyed out low level metrics are not optimized for in these experiments.
Low-Level High-Level
Subject 1 FT 1HR Baseline E1 0.031 0.234 58.31% 62.56% 58.51% 58.71% 0.947 0.611
Subject 1 FT 1HR AAMAX E1 0.095 0.254 77.20% 85.98% 78.80% 79.39% 0.833 0.503
Subject 1 FT 1HR Baseline E160 0.094 0.239 71.54% 80.10% 73.11% 74.02% 0.875 0.537
Subject 1 FT 1HR AAMAX E160 0.105 0.259 76.77% 86.20% 80.08% 80.14% 0.828 0.496
Subject 1 FT 2HR Baseline E1 0.049 0.242 62.41% 68.06% 62.64% 64.80% 0.925 0.583
Subject 1 FT 2HR AAMAX E1 0.112 0.275 79.46% 88.29% 83.28% 82.61% 0.803 0.485
Subject 1 FT 2HR Baseline E160 0.110 0.270 75.32% 84.62% 76.66% 78.55% 0.845 0.516
Subject 1 FT 1HR AAMAX E160 0.114 0.285 80.00% 88.50% 83.19% 83.05% 0.799 0.480
Subject 1 FT 4HR Baseline E1 0.079 0.241 65.84% 73.04% 64.19% 68.10% 0.914 0.574
Subject 1 FT 4HR AAMAX E1 0.129 0.265 80.83% 90.93% 86.66% 86.17% 0.767 0.460
Subject 1 FT 4HR Baseline E160 0.124 0.253 77.86% 89.18% 83.10% 83.26% 0.801 0.480
Subject 1 FT 4HR AAMAX E160 0.127 0.273 81.73% 91.29% 86.23% 86.26% 0.763 0.453
Figure 7: Comparing finetuning reconstruction performance over epochs with limited data. From left to right: (a) Finetuning Subject 1 with 250 images or 1 hour of data (b) Finetuning Subject 1 with 500 images or 2 hours of data (c) Finetuning Subject 1 with 1000 images or 3 hours of data. AAMax not only outperforms normal end-to-end training, but the performance at Epoch 1 is better than the baseline’s performance at Epoch 160.
Appendix G Ablation Study on the Two Phases of Adapter Alignment Fine-Tuning
We analyze the role of Step 1 (adapter alignment) relative to full end-to-end fine-tuning. In this experiment, we fine-tuned Subject 2 with 1 hour of data using Subject 1 as the reference. Figure 8 illustrates the reconstruction performance across epochs for both phases.
During Step 1, where only the adapter is optimized, performance improves steadily as the adapter overfits to the shared images. This phase alone provides a substantial boost over the unaligned baseline, and after sufficient epochs the adapter can achieve strong alignment across subjects. However, the improvement plateaus, as Step 1 only modifies the adapter and does not allow deeper components of the pipeline to adapt.
Introducing end-to-end fine-tuning after adapter alignment leads to an immediate jump in performance. Even after a single epoch of end-to-end updates, the model surpasses what hundreds of adapter-only epochs can achieve. We stop at around 200 epochs as the relative gain plateaus again. While the improvement in performance by a few percent from end-to-end fine-tuning might not seem significant when compared to the gain from Step 1, it can make the difference between coherent images and gibberish when working with limited data and when reconstruction fidelity is low.
Figure 8: Reconstruction performance across epochs for Subject 2 fine-tuned with 1 hour of data using Subject 1 as the reference. The shaded regions denote Step 1 (adapter alignment) and Step 2 (end-to-end fine-tuning). Performance steadily increases during adapter alignment but plateaus short of optimal results. Transitioning to end-to-end fine-tuning yields an immediate jump in accuracy and ultimately outperforms adapter-only training. The dashed lines indicate the transition from AA to E2E. Performance at Epoch 300 of AA is the same as Epoch 0 of E2E, since no further fine-tuning has been done yet.
Appendix H Additional Scaling Analysis
The architectural setup used by MindEye2 leverages a 4096-dimensional shared space with BigCLIP and Stable Diffusion XL (SDXL), requiring substantially greater computational resources than our main experiments, which employ Versatile Diffusion. To evaluate whether the benefits of Adapter Alignment (AAMax) persist across different model capacities, we conducted a scaling law analysis varying both the dimensionality of the shared space and the choice of generative backbone.
H.1 Empirical Observations
@@
These findings indicate that AAMax is robust to architectural choice and continues to provide benefits regardless of the CLIP encoder or diffusion model employed.
H.2 Projection to 4096 Dimensions with Stable Diffusion XL
The key insight from these experiments is that adapter alignment is a training paradigm, not an architectural modification. Our results demonstrate that AA is agnostic to the representational scale and generative backbone, and scales effectively with larger models. While large-scale pipelines like MindEye2 set upper bounds on reconstruction quality under high compute budgets, adapter alignment provides a complementary framework that improves efficiency and accessibility. Importantly, its benefits persist even when extended to high-capacity architectures, making it a broadly applicable strategy for fMRI-based reconstruction.